Exercise 15-18 (Part Level Submission) Cheyenne Company reported the following amounts in the stockholders equity section of its December 31, 2016, balance sheet. Preferred stock, 1096, $100 par (10,000 shares authorized, 1,800 shares issued) $180,000 Common stock, $5 par (103,500 shares authonized, 20,700 shares issued) Additional paid-in capital Retained earnings 103,500 114,000 423,000 $820,500 Total During 2017, Cheyenne took part in the following transactions concerning stockholders equity 1. Paid the annual 2016 $10 per share dividend on preferred stock and a $2 per share dividend on common stock. These dividends had been declared on December 31, 2016. 2. Purchased 1,500 shares of its own outstanding common stock for $43 per share. Cheyenne uses the cost method. 3. Reissued 600 treasury shares for land valued at $27,100. 4. Issued 550 shares of preferred stock at $105 per share. 5. Declared a 10% stock dividend on the outstanding common stock when the stock is selling for $41 per share. 6. Issued the stock dividend. 7. Declared the annual 2017 $10 per share dividend on preferred stock and the $2 per share dividend on common stock. These dividends are payable in 2018. ▼ (a) Prepare journal entries to record the transactions described above. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is reauired, selert Nn Entru for the arcount titles and enter for the amounte.)
Answer #1
Cheyenne Company
S.No. General, Journal Debit Credit
1) Dividend Payable-Common Stock(20700*$2) $ 41,400.00
Dividend Payable-Preferred Stock(1800*$10) $ 18,000.00
   To Cash $ 59,400.00
(Being paid annual dividend on 1800 preference shares @$10 per share and 20700 Equity shares @$2 per share)
2) Treasury Stock(1500*$43) $ 64,500.00
To Cash $ 64,500.00
(Purchased 1500 own stock @$43)
3) Land $ 27,100.00
   To Treasury Stock(600 *$43) $ 25,800.00
   To Paid in capital for treasury stock $   1,300.00
(Being reissued 600 shares for land)
4) Cash(550*$105) $ 57,750.00
    To Preferred Stock(550*$100) $ 55,000.00
     To Paid in Capital excess of par-Preferred stock $   2,750.00
(Issued 550 Preferred shares @$105)
5) Retained Earnings(1980*$41) $ 81,180.00
*    To Common Stock Dividend Distributable(1980*$5) $   9,900.00
To Paid in Capital excess of par -common stock=1980*($41-$5) $ 71,280.00
(Being declared a 10% stock dividend on the outstanding common stock)
* Common Stock 20700
Less: Treasury Stock -1500
Add: Reissued for land 600
19800
(20700-1500+600)=19800*10%=1980
6) Common Stock dividend distributable $   9,900.00
    To Common Stock $   9,900.00
(Being amount of dividend distributed)
7) Retained Earnings $ 67,060.00
**    To Dividend Payable-Preferred Stock(1800+550)*$10 $ 23,500.00
***    To Dividend Payable-Common Stock=(19800+1980)*$2=(21780*$2) $ 43,560.00
(Being amount of Dividend declared on 2350 @$10 per share on preferred stock,21780 Equity shares @$2 per share on common stock)
Preferred Stock available at the time of declared the annual 2017 dividend
Preferred stock 1800
Add: New Preferred stock issued 550
Preferred Stock available 2350
Common Stock available at the time of declared the annual 2017 dividend
Common Stock 20700
Less: Treasury Stock -1500
Add: Reissued for land 600
Stock dividend 1980
21780